what is the area under a curve with the equation y= negative 2x to the 2nd power + 6x + 36
Answer by Alan3354(69443)

what is the area under a curve with the equation y= negative 2x to the 2nd power + 6x + 36
------------
y = -2x^2 + 6x + 36
Assuming you mean bounded by the x-axis:
It intersects the x-axis at x = -3 and x = +6
-------------
INT(x) = -2x^3/3 + 3x^2 + 36x (Ignore the constant of integration)
INT(6) = -144 + 108 + 216 = 180
INT(-3) = 18 + 27 - 108 = -63
-----
INT(6) - INT(-3) = 243 sq units
Area = 243 sq units